On January 1, 2017, Boston Enterprises issues bonds that have a $3,400,000 par value, mature in 20 years, and pay 9% interest semiannually on June 30 and December 31. The bonds are sold at par.

1. How much interest will Boston pay (in cash) to the bondholders every six months?
2. Prepare journal entries to record (a) the issuance of bonds on January 1, 2017; (b) the first interest payment on June 30, 2017; and (c) the second interest payment on December 31, 2017.
3. Prepare the journal entry for issuance assuming the bonds are issued at (a) 98 and (b) 102.
Answer #1
Concepts and reason

Accounting: Accounting is a process of recording the transactions, classifying them in a specific manner, and then the process of summarizing and analyzing is carried out to interpret the results. It is a process of preserving the accounts.

Financial accounting is a process of preparing reports to provide all financial information to both the internal and external users of an organization. The financial statements that are prepared under the financial accounting are examined by independent certified public accountants, at the year-end, who would express their opinion on the fairness of the reports shown by a company.

Transaction: Transaction is an act of buying or selling goods or rendering any service that is reliably measured in terms of money.

Journal entry: Journal entry is the recording of transactions in a systematic manner as they occur. Thus, it is a summary of all the transactions which has debit and credit aspects recorded chronologically.

Fundamentals

Bonds: Bonds are promises that are made between two or more persons legally. Bonds are issued by the companies for raising their capital balance. Interest is received at regular intervals by a purchaser.

Bonds payable: Bonds payable is a long-term legal agreement issued by a company to several bondholders. Bondholders are entitled to receive a specified interest on the bonds bought by them. Bonds payable is a long-term liability, and therefore reported under long-term liabilities section in the balance sheet.

Bonds issued at discount: When the selling price of the bond is less than the face value represented in the bond, then the bond is said to be issued at discount.

Bonds issued at premium: When the selling price of the bond is higher than the face value represented in the bond, then the bond is said to be issued at premium.

Interest expense: This is a non-operating expense, and it will be shown on the income statement. Generally, this expense is considered as an amount incurred by the company for using the other resources like borrowing loan from bank or purchasing a note and so on.

Golden rules of accounting:

Accounts
Personal
Real
Debit
The receiver
What comes in
All expenses and
losses
Credit
The giver
What goes out
All income and

Rules for debit and credit:

When assets increase, debit them; when assets decrease, credit them.

When liabilities increase, credit them; when liabilities decrease, debit them.

When stockholders’ equity increases, credit it; when stockholders’ equity decreases, debit it.

When expenses and losses increase, debit them; when expenses and losses decrease, credit them.

When incomes and gains increase, credit them; when incomes and gains decrease, debit them.

1)

Calculate the interest paid by “B” to the bondholders:

Interestpayment=(Parvalueofbond×Semiannualinterestrate)=$3,400,000×4.5%=$153,000\begin{array}{c}\\{\rm{Interest payment = }}\left( \begin{array}{l}\\{\rm{Par value of bond}}\\\\ \times {\rm{Semi annual interest rate}}\\\end{array} \right)\\\\{\rm{ = \$ 3,400,000}} \times {\rm{4}}{\rm{.5\% }}\\\\{\rm{ = \$ 153,000}}\\\end{array}

Therefore, the interest paid by “B” to the bondholders for every six months is $153,000.

Working notes:

Calculate the semi-annual interest rate:

Semiannualinterestrate}=Interestrate2=9%2=4.50%\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Semiannual interest }}\\\\{\rm{rate}}\\\end{array} \right\}{\rm{ = }}\frac{{{\rm{ Interest rate}}}}{2}\\\\ = \frac{{9\% }}{2}\\\\ = 4.50\% \\\end{array}

Therefore, the semi-annual interest rate is 4.50%.

2.a)

Prepare the journal entry to record the issuance of bonds on January 1, 2017:

Credit
Debit
$3.400.000
Date Accounts title and explanation
1-Jan-17 Cash
Bonds payable
To record the issuance of bond at par

Therefore, the cash account is debited and the bonds payable is credited for $3,400,000.

2.b)

Prepare the journal entry to record the issuance of bond at par:

Credit
Debit
$153,000
Date Accounts title and explanation
1-Jun-17 Bonds interest expense
Cash
(To record the first interest

Therefore, the bonds interest expense is debited and the cash account is credited for $153,000.

2.c)

Prepare the journal entry to record the second interest payment on December 31, 2017:

Date Accounts title and explanation
31-Dec-17 Bonds interest expense
Cash
(To record the second interest payment)
Debit Credi

Therefore, the bonds’ interest expense is debited and the cash account is credited for $153,000.

3.a)

Prepare the journal entry to record bond issued at $98:

Credit
Debit
$3,332,000
$68,000
Date Accounts title and explanation
Cash
Discount on bonds payable
Bonds payable
(To record t

Therefore, the cash account is debited for $3,332,000, discount on bonds payable is $68,000, and bonds payable is credited for 3,400,000.

Working notes:

Calculate the issue of bonds for $98.

Bondsissuedforcash=Priceofbond×Parvalueofthebond=0.98×$3,400,000=$3,468,000\begin{array}{c}\\{\rm{Bonds issued for cash = Price of bond}} \times {\rm{Par value of the bond}}\\\\{\rm{ = 0}}{\rm{.98}} \times {\rm{\$ 3,400,000}}\\\\{\rm{ = \$ 3,468,000}}\\\end{array}

Therefore, the bonds issued for cash is $3,468,000.

Calculate the discount on bonds payable:

Discountonbondspayable=(ParvalueofthebondBondsissuedforcash)=$3,400,000$3,332,000=$68,000\begin{array}{c}\\{\rm{Discount on bonds payable = }}\left( \begin{array}{l}\\{\rm{ Par value of the bond}}\\\\ - {\rm{Bonds issued for cash}}\\\end{array} \right)\\\\ = \$ 3,400,000 - \$ 3,332,000\\\\ = \$ 68,000\\\end{array}

Therefore, the discount on bonds payable is $68,000.

3.b)

Prepare the journal entry to record bond issued at $102:

Credit
Debit
$3,468,000
Date Accounts title and explanation
Cash
Premium on bonds payable
Bonds payable
(To record the issue

Therefore, the cash account is debited for $3,468,000, premium on bonds payable is credited for $68,000, and bonds payable is credited for 3,400,000.

Working notes:

Calculate the issue of bonds for $102.

Bondsissuedforcash=Priceofbond×Parvalueofthebond=1.02×$3,400,000=$3,468,000\begin{array}{c}\\{\rm{Bonds issued for cash = Price of bond}} \times {\rm{Par value of the bond}}\\\\{\rm{ = 1}}{\rm{.02}} \times {\rm{\$ 3,400,000}}\\\\{\rm{ = \$ 3,468,000}}\\\end{array}

Therefore, the bonds issued for cash is $3,468,000.

Calculate the discount on bonds payable:

Discountonbondspayable=(ParvalueofthebondBondsissuedforcash)=$3,468,000$3,400,000=$68,000\begin{array}{c}\\{\rm{Discount on bonds payable = }}\left( \begin{array}{l}\\{\rm{ Par value of the bond}}\\\\ - {\rm{Bonds issued for cash}}\\\end{array} \right)\\\\ = \$ 3,468,000 - \$ 3,400,000\\\\ = \$ 68,000\\\end{array}

Therefore, the discount on bonds payable is $68,000.

Ans: Part 1

The interest paid by “B” to the bondholders for every six months is $153,000.

Part 2.a

Date Accounts title and explanation
1-Jan-17 Cash
Bonds payable
(To record the issuance of bond at par)
Debit Credit
$3,400,0

Part 2.b

Credit
Debit
$153,000
Date Accounts title and explanation
1-Jun-17 Bonds interest expense
Cash
(To record the first interest

Part 2.c

Credit
Debit
$153,000
| Date Accounts title and explanation
31-Dec-17 Bonds interest expense
Cash
(To record the second inter

Part 3.a

Credit
Debit
$3,332,000
$68,000
Date Accounts title and explanation
Cash
Discount on bonds payable
Bonds payable
(To record t

Part 3.b

Credit
Debit
$3,468,000
Date Accounts title and explanation
Cash
Premium on bonds payable
Bonds payable
(To record the issue